OpenStudy (anonymous):

which of the following is equal to square root 18x^7y^6

OpenStudy (anonymous):

i know the answer is 3x^3y^3sqrt2 but how would i figure this out

OpenStudy (anonymous):

Here's the first step: \[\sqrt{2*3^2*x^6xy^6}\]

OpenStudy (anonymous):

Second step: Separate terms you can take the sqrt of... \[\sqrt{3^2x^6y^6}*\sqrt{2x}\] 3rd Step: \[3x^3y^3\sqrt{2x}\]

OpenStudy (anonymous):

is that any clearer? or do you need me to explain any of the steps?

OpenStudy (anonymous):